OpenStudy (anonymous):

Write the equation for a circle with its center at (2,-5) and a radius of 6.

OpenStudy (cwrw238):

the general form of the equation is (x-a)^2 + (y+b)^2 = r^2 where center is (a,b) and radius is r

OpenStudy (anonymous):

\[(x-3)^2 + (y+6)= \]

OpenStudy (cwrw238):

* sorry thats (y-b)^2

OpenStudy (cwrw238):

(x - 2)^2 + (y + 5)^2 = 36

OpenStudy (anonymous):

Is any one good at marine biology? cuz I need help plz. Easy medal!